A TV manufacturer claims that the life span of its regular TV sets is longer than 10 years. Using a random sample of their 25 TV sets, the average life span is found to be 11.9 years with a standard deviation of 1.8 years. Test the hypothesis that the TV sets' life span is longer than 10 years at cx = 0.10


1
Expert's answer
2022-06-16T05:58:07-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=10"

"H_1:\mu>10"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=24" and the critical value for a right-tailed test is "t_c =1.317836."

The rejection region for this right-tailed test is "R = \{t:t>1.317836\}."

The t-statistic is computed as follows:


"t=\dfrac{11.9-10}{1.8/\sqrt{25}}=\dfrac{110-100}{10/\sqrt{16}}=5.2778"


Since it is observed that "t=5.2778>1.317836=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=24" degrees of freedom, "t=5.2778" is "p=0.00001," and since "p=0.00001<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 10, at the "\alpha = 0.10" significance level.